OpenStudy (nsh4267):

Find a general form of an equation for the perpendicular bisector of the segment AB. A(4, 2), B(−2, 10)

OpenStudy (anonymous):

find the line going through A and B by first finding its gradient (m=\[m=(y2-y1)\div(x2-x1)\], then find the point which is in the middle of AB by using the equation for the median point (x=(x2+x1)/2, y=(y2+y1)/2 then, knowing that the perpendicular line will go through this median point, and knowing that the gradient of the perpendicular line is opposite and inverse to the gradient of the line passing through AB, find the equation of the line

OpenStudy (nsh4267):

which is 1,6

OpenStudy (nsh4267):

thats the midpoint

OpenStudy (nsh4267):

the slope is 3/4

OpenStudy (anonymous):

ok, so just express this as a general equation